# Subsampled Ensemble Can Improve Generalization Tail Exponentially

### v3

@@ 2.2 A more general procedure
We next present a more general procedure called $\mathsf{ROVE}$ that applies to continuous space where simple majority vote in Algorithm 1 can lead to degeneracy, i.e., all learned models appear exactly once in the pool. Moreover, this general procedure relaxes our dependence on $|\Theta|$ in the bound in Theorem 1.
@@
$\mathsf{ROVE}$, displayed in Algorithm 2, proceeds initially the same as $\mathsf{MoVE}$ in repeatedly subsampling data and training the model using $\mathcal{A}$. However, in the aggregation step, instead of using a simple majority vote, $\mathsf{ROVE}$ outputs, among all the trained models, the one that has the highest likelihood of being $\epsilon$-optimal. This $\epsilon$-optimality avoids the degeneracy of the majority vote and, moreover, since we have restricted our output to the collection of trained models, the corresponding likelihood maximization is readily doable by simple enumeration. In addition, it helps reduce competition for votes among the best models as each subsample can now vote for multiple candidates, ensuring a high vote count for each of the top models even when there are many of them. This makes $\mathsf{ROVE}$ more effective than $\mathsf{MoVE}$ in the case of multiple (near) optima as our experiments will show. We have the following theoretical guarantees for Algorithm 2:
Theorem 2 (Finite-sample bound for Algorithm 2)
@@
Recall the tail $\mathcal{E}_{k,\delta}$ of the base excess risk from Theorem 1. Consider Algorithm 2 with data splitting, i.e., $\mathsf{ROVEs}$. Let $T_{k}(\cdot):=\mathbb{P}(\sup_{\theta\in\Theta}\lvert(1/k)\sum_{i=1}^{k}l(\theta,z_{i})-L(\theta)\rvert>\cdot)$ be the tail function of the maximum deviation of the empirical objective estimate. For every $\delta>0$, if $\epsilon$ is chosen such that $\mathbb{P}\left(\epsilon\in[\underline{\epsilon},\overline{\epsilon}]\right)=1$ for some $0<\underline{\epsilon}\leq\overline{\epsilon}<\delta$ and $T_{k_{2}}\left((\delta-\overline{\epsilon})/2\right)+T_{k_{2}}\left(\underline{\epsilon}/2\right)<1/5$, then
$\mathbb{P}\Big{(}L(\hat{\theta}_{n})>\min_{\theta\in\Theta}L(\theta)+2\delta\Big{)}\leq$ $B_{1}\left[3\min\left(e^{-2/5},C_{1}T_{k_{2}}\left(\dfrac{\min(\underline{\epsilon},\delta-\overline{\epsilon})}{2}\right)\right)^{\frac{n}{2C_{2}k_{2}}}+e^{-{B_{2}}/{C_{3}}}\right]$ (10)
$+$ $\min\left(e^{-(1-\mathcal{E}_{k_{1},\delta})/C_{4}},C_{5}\mathcal{E}_{k_{1},\delta}\right)^{\frac{n}{2C_{6}k_{1}}}+e^{-B_{1}(1-\mathcal{E}_{k_{1},\delta})/C_{7}},$
where $C_{1},C_{2},C_{3}$ are the same as those in Theorem 1, and $C_{4},C_{5},C_{6},C_{7}$ are universal constants.
@@
Theorem 2 provides an exponential excess risk tail, regardless of discrete or continuous space. The first line in the bound (10) is inherited from the bound (9) for $\mathsf{MoVE}$ from majority to $\epsilon$-optimality vote. In particular, the multiplier $\left|\Theta\right|$ in (9) is now replaced by $B_{1}$, the number of retrieved models. The second line in (10) bounds the performance sacrifice due to the restriction to Phase I model candidates.
@@
$\mathsf{ROVE}$ may be carried out with the data split between the two phases, in which case it’s referred to as $\mathsf{ROVEs}$. Data splitting makes the procedure theoretically more tractable by avoiding inter-dependency between the phases but sacrifices some statistical power from halving the data size. Empirically we find $\mathsf{ROVE}$ to be overall more effective.
@@
The optimality threshold $\epsilon$ is allowed to be chosen in a data-driven way and the main goal guiding this choice is to be able to distinguish models of different qualities. In other words, $\epsilon$ should be chosen to create enough variability in the likelihood of being $\epsilon$-optimal across models. In our experiments, we find it a good strategy to choose an $\epsilon$ that leads to a maximum likelihood around $1/2$.
@@
Lastly, our main theoretical results, Theorems 1 and 2, are derived using several novel techniques. First, we develop a sharper concentration result for U-statistics with binary kernels, improving upon standard Bernstein-type inequalities (e.g., Arcones (1995); Peel et al. (2010)). This refinement ensures the correct order of the bound, particularly (9), which captures the convergence of both the bootstrap approximation and the base learner, offering insights into the robustness of our methods for fast-converging base learners. Second, we perform a sensitivity analysis on the regret for the original problem (1) relative to the surrogate optimization (4), translating the superior generalization in the surrogate problem into accelerated convergence for the original. Finally, to establish asymptotic consistency for Algorithm 2 without data splitting, we develop a uniform law of large numbers (LLN) for the class of events of being $\epsilon$-optimal, using direct analysis of the second moment of the maximum deviation. Uniform LLNs are particularly challenging here because, unlike fixed classes in standard settings, this class dynamically depends on subsample size $k_{2}$ as $n\to\infty$.
3 Numerical experiments
In this section, we numerically test Algorithm 1 ($\mathsf{MoVE}$), Algorithm 2 with ($\mathsf{ROVEs}$) and without ($\mathsf{ROVE}$) data splitting in training neural networks for regression problems and solving stochastic programs. Additional experimental results are provided in Appendix D due to space constraints. The code is available at: https://anonymous.4open.science/r/vote_ensemble.
To empirically determine well-performing configurations for general use, we performed a comprehensive hyperparameter profiling of our algorithms in Appendix D.3. Below, we summarize the recommended configurations used in all experiments presented in this section (except Figure 4): 1) For discrete space $\Theta$, use $k=\max(10,n/200),B=200$ for $\mathsf{MoVE}$, and $k_{1}=k_{2}=\max(10,n/200),B_{1}=20,B_{2}=200$ for $\mathsf{ROVE}$ and $\mathsf{ROVEs}$; 2) For continuous space $\Theta$, use $k_{1}=\max(30,n/2),k_{2}=\max(30,n/200),B_{1}=50,B_{2}=200$ for $\mathsf{ROVE}$ and $\mathsf{ROVEs}$; 3) The $\epsilon$ in $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ is selected such that $\max_{\theta\in\mathcal{S}}(1/B_{2})\sum_{b=1}^{B_{2}}\mathbbm{1}(\theta\in\widehat{\Theta}^{\epsilon,b}_{k_{2}})\approx 1/2$.
3.1 Neural networks for regression
We consider regression problems with multilayer perceptrons (MLPs) on both synthetic and real data. The base learning algorithm splits the data into training (70%) and validation (30%), and uses Adam to minimize mean squared error (MSE), with early stopping triggered when the validation improvement falls below 3% between epochs. The architecture details of the MLPs are provided in Appendix D.1. Note that $\mathsf{MoVE}$ is not included in this comparison as it’s applicable to discrete problems only.
Setup for Synthetic Data
Input-output pairs $(X,Y)$ are generated as $Y=(1/50)\cdot\sum_{j=1}^{50}\log(X_{j}+1)+\varepsilon$, where each $X_{j}$ is drawn independently from $\mathrm{Unif}(0,2+198(j-1)/49)$, and the noise $\varepsilon$ is independent of $X$ with zero mean. We consider both standard Gaussian noise and Pareto noise $\varepsilon=\varepsilon_{1}-\varepsilon_{2}$, where each $\varepsilon_{i}\sim\mathrm{Pareto}(2.1)$. The out-of-sample performance is estimated on a common test set of one million samples. Each algorithm is repeatedly applied to $200$ independently generated datasets to assess the average and tail performance.
Setup for Real Data
We use six datasets from the UCI Machine Learning Repository (Blake, 1998): Wine Quality (Cortez et al., 2009), Bike Sharing (Fanaee-T, 2013), Online News (Fernandes et al., 2015), Appliances Energy (Candanedo, 2017), Superconductivity (Hamidieh, 2018), and Gas Turbine Emission (gas, 2019). Each dataset is standardized (zero mean, unit variance). To evaluate the average and tail performance, we permute each dataset $100$ times, and each time use the first half for training and the second for testing.
@@ 3.1 Neural networks for regression
Result.
@@
As shown in Figure 1, in heavy-tailed noise settings (Figures 1a–1c), both $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ significantly outperform the base algorithm in terms of both expected out-of-sample MSE and tail performance under all sample sizes $n$. Notably, the performance improvement becomes more pronounced with deeper networks ($H=8$), indicating that the benefits of $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ are more apparent in models with higher expressiveness and lower bias.
@@
In light-tailed settings (Figures 1d–1f), $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ show comparable expected out-of-sample performance to the base when $H=4$, but outperform the base as $H$ increases. Additionally, $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ outperform the base in tail probabilities even when $H=4$. This indicates that $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ provide better generalization as the model complexity grows even for light-tailed problems. Similar results for MLPs with $2$ and $6$ hidden layers can be found in Appendix D.4, where results on least squares regression and Ridge regression are also provided.
@@
On real datasets (Figure 2), $\mathsf{ROVE}$ exhibits much lighter tails compared to the base on three out of six datasets, and similar tail behavior on the other three. $\mathsf{ROVEs}$, however, underperforms the base in these real-world scenarios, potentially due to the data split that compromises its statistical power.
@@ 3.2 Stochastic programs
We consider four discrete stochastic programs: resource allocation, supply chain network design, maximum weight matching, and stochastic linear programming, alongside one continuous mean-variance portfolio optimization. All problems are designed to possess heavy-tailed uncertainties. For the stochastic linear program, instances with varying tail heaviness are explored to study its impact on algorithm performance. The base learning algorithm for all the problems is the SAA. Detailed descriptions of the problems are deferred to Appendix D.2 and results using DRO as the base algorithm are provided in Appendix D.4.
(a) Resource allocation.
(b) Network design.
(c) Portfolio optimization.
(d) Maximum weight matching.
(e) Linear program (multiple optima).
(f) Tail of portfolio opt., $n=2^{16}$.
Figure 3: Results for stochastic programs. (a)-(e): Expected out-of-sample costs with $95\%$ confidence intervals. (f): Tail probabilities of out-of-sample costs for mean-variance portfolio optimization. All maximization problems are converted to minimization by negating their objectives, and the generic term “cost” refers to the minimizing objective.
(a) Influence of tail heaviness.
(b) Running time comparison.
Figure 4: (a): Influence of tail heaviness in the stochastic linear program with multiple optima with $n=10^{6}$. Hyperparameters: $k=50,B=2000$ for $\mathsf{MoVE}$, $k_{1}=k_{2}=50,B_{1}=200,B_{2}=5000$ for $\mathsf{ROVE}$ and $\mathsf{ROVEs}$. The tail heaviness parameter corresponds to the mean of the Pareto random coefficient. (b): Running time for supply chain network design. Hyperparameters: $k=10,B=200$ for $\mathsf{MoVE}$, $k_{1}=k_{2}=10,B_{1}=20,B_{2}=200$ for $\mathsf{ROVE}$ and $\mathsf{ROVEs}$. “Sequential” refers to sequential processing of the subsamples; “Parallel” refers to parallel processing with 8 CPU cores.
Result.
@@
Figure 3 shows that our ensembling methods generally outperform the base algorithm in all cases, except for the linear program case (Figure 3(e)). Notably, $\mathsf{ROVE}$ still outperforms the base in the linear program case, demonstrating its robustness, while $\mathsf{MoVE}$ performs slightly worse than the base under small sample sizes. Comparing $\mathsf{ROVE}$ and $\mathsf{ROVEs}$, $\mathsf{ROVE}$ consistently exhibits superior performance than $\mathsf{ROVEs}$ in all cases.
@@
When there is a unique optimal solution, $\mathsf{MoVE}$ and $\mathsf{ROVE}$ perform similarly, both generally better than $\mathsf{ROVEs}$, as seen in Figures 3(a)-3(d). However, in cases with multiple optima (Figures 3(e) and 4(a)), the performance of $\mathsf{MoVE}$ deteriorates while $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ stay strong. This is in accordance with our discussion on the advantage of $\epsilon$-optimality vote in Section 2.2. Additional results in Appendix D.4 shall further explain that optima multiplicity weakens the base learner for $\mathsf{MoVE}$ in the sense of decreasing the $\eta_{k,\delta}$ and hence inflating the tail bound in Theorem 1.
As shown in Figure 4(a), the performance gap between $\mathsf{ROVE}$, $\mathsf{ROVEs}$, and the base algorithm becomes increasingly significant as the tail of the uncertainty becomes heavier. This supports the effectiveness of $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ in handling heavy-tailed uncertainty, where the base algorithm’s performance suffers. Note that here $\mathsf{MoVE}$ behaves similarly as the base due to optima multiplicity.
@@
The running time comparison in Figure 4(b) shows that, despite requiring multiple runs on subsamples, our ensembling methods do not introduce a significantly higher computational burden compared to running the base algorithm on the full sample, and can even be advantageous under large sample sizes. This is because, in problems like DRO (Ben-Tal et al., 2013; Mohajerin Esfahani & Kuhn, 2018) and two-stage stochastic programming, solving the optimization on the full sample often leads to a substantial increase in problem size, as the decision space and constraints grow at least linearly with the sample size. Subsampled optimizations, as performed in our approach, result in smaller, more manageable problems that can be solved more efficiently. Moreover, our theory indicates that solving more than $\mathcal{O}(n/k)$ subsamples does not further improve generalization performance, ensuring that computational efficiency is maintained. Additionally, parallel processing of subsamples further reduces computational time.
@@
Finally, among the three proposed ensemble methods, $\mathsf{ROVE}$ is the preferred choice over $\mathsf{MoVE}$ and $\mathsf{ROVEs}$ for general use as it’s applicable to both discrete and continuous problems and consistently delivers superior and stable performance across all scenarios.

### v4

@@ 2.2 A More General Procedure
We next present a more general procedure called $\mathsf{ROVE}$ that applies to continuous space where the simple majority vote in $\mathsf{MoVE}$ can lead to degeneracy, i.e., all learned models appear exactly once in the pool. Moreover, this general procedure relaxes the dependence on $|\Theta|$ in the bound (9).
@@
$\mathsf{ROVE}$, displayed in Algorithm 2, proceeds initially the same as $\mathsf{MoVE}$ in repeatedly subsampling data and training the model using $\mathcal{A}$. However, in the aggregation step, instead of using a simple majority vote, $\mathsf{ROVE}$ outputs, among all the trained models, the one that has the highest likelihood of being $\epsilon$-optimal. This $\epsilon$-optimality avoids the degeneracy of the majority vote. Moreover, since we have restricted our output to the collection of retrieved models, the corresponding likelihood maximization is readily doable by simple enumeration. In addition, it helps reduce competition for votes among best models, as each subsample can now vote for multiple candidates, ensuring a high vote count for each of the top models even when there are many of them. This makes $\mathsf{ROVE}$ more effective than $\mathsf{MoVE}$ in the case of multiple (near) optima as our experiments will show. We have the following theoretical guarantees for Algorithm 2.
Theorem 2.3 (Informal bound for Algorithm 2).
@@
Let $\mathcal{E}_{k,\delta}$ be the excess risk tail defined in Theorem 2.1. Consider Algorithm 2 with data splitting, i.e., $\mathsf{ROVEs}$. Let $T_{k}(\cdot):=\mathbb{P}(\sup_{\theta\in\Theta}\lvert(1/k)\sum_{i=1}^{k}l(\theta,z_{i})-L(\theta)\rvert>\cdot)$ be the tail function of the maximum deviation of the empirical objective estimate. Then, for every $\delta>0$, under mild conditions on $\epsilon$ and $T_{k_{2}}(\cdot)$, it holds that
$\mathbb{P}$ $\Big{(}L(\hat{\theta}_{n})>\min_{\theta\in\Theta}L(\theta)+2\delta\Big{)}\leq B_{1}\Big{[}3\exp(-\Omega(n/k_{2}))$ (11)
$+\exp(-\Omega(B_{2}))\Big{]}+\exp(-\Omega(n/k_{1}))+\exp(-\Omega(B_{1})),$
where $\Omega(\cdot)$ contains multiplicative coefficients depending on $T_{k_{2}}(\cdot)$, $\epsilon$, $\delta$ and $\mathcal{E}_{k_{1},\delta}$.
Consider Algorithm 2 without data splitting, i.e., $\mathsf{ROVE}$, and discrete space $\Theta$. Assume $\lim_{k\to\infty}T_{k}(\delta)=0$ for all $\delta>0$. Then, for every fixed $\delta>0$, under mild conditions, it holds that $\lim_{n\to\infty}\mathbb{P}(L(\hat{\theta}_{n})>\min_{\theta\in\Theta}L(\theta)+2\delta)\to 0$.
@@
The formal statement of Theorem 2.3 can be found in Appendix C.4. Theorem 2.3 provides an exponential excess risk tail, regardless of discrete or continuous space. The terms in the square bracket of (11) are inherited from the bound (10) for $\mathsf{MoVE}$ with the majority vote replaced by $\epsilon$-optimality vote. In particular, the multiplier $\left|\Theta\right|$ in (10) is now replaced by $B_{1}$, the number of retrieved models from Phase I. The last two terms in (11) bound the performance sacrifice due to the restriction to the retrieved models.
@@
$\mathsf{ROVE}$ may be carried out with the data split between the two phases, in which case it’s referred to as $\mathsf{ROVEs}$. Data splitting makes the procedure theoretically more tractable by avoiding inter-dependency between the phases but sacrifices some statistical power by halving the data size. Empirically we find it more effective not to split data.
@@
The optimality threshold $\epsilon$ is allowed to be chosen in a data-driven way and the main goal guiding this choice is to distinguish models of different qualities. In other words, $\epsilon$ should be chosen to create enough variability in the likelihood of being $\epsilon$-optimal across models. In our experiments, we find it a good strategy to choose an $\epsilon$ that leads to a maximum likelihood around $1/2$.
@@
Technical Novelty.    Our main theoretical results, Theorems 2.1 and 2.3, are derived using several novel techniques. First, we develop a sharper concentration result for U-statistics with binary kernels, improving upon standard Bernstein-type inequalities (e.g., (Arcones, 1995; Peel et al., 2010)). This refinement ensures the correct order of the bound, particularly (10), which captures the convergence of both the bootstrap approximation and the base learner, offering insights into the robustness of our methods for fast-converging base learners. Second, we perform a sensitivity analysis on the regret for the original problem (1) relative to the surrogate optimization (5), translating the superior generalization in the surrogate problem into accelerated convergence for the original. Finally, to establish asymptotic consistency for Algorithm 2 without data splitting, we develop a uniform law of large numbers (LLN) for the class of events of being $\epsilon$-optimal, using direct analysis of the second moment of the maximum deviation. Uniform LLNs are particularly challenging here because, unlike fixed classes in standard settings, this class dynamically changes with the subsample size $k_{2}$ as $n\to\infty$.
@@
(a) Pareto shape $=2.1$.
(b) Pareto shape $=1.5$.
(c) Pareto shape $=1.1$.
Figure 2: Comparison with bagging on synthetic data in terms of expected out-of-sample costs (MSE) with $95\%$ confidence intervals under varying degrees of tail heaviness. Hyperparameters: $k_{1}=\max(30,n/2),k_{2}=\max(30,n/1000),B_{1}=50,B_{2}=1000$.
@@
(a) Resource allocation.
(b) Portfolio optimization.
(c) Maximum weight matching.
(d) Linear program (multiple optima).
(e) Network design.
(f) Network design running time.
Figure 4: Results for stochastic programs. (a)-(e): Expected out-of-sample costs with $95\%$ confidence intervals. (f): Running time comparison in the network design problem.
3 Numerical Experiments
In this section, we numerically test Algorithm 1 ($\mathsf{MoVE}$), Algorithm 2 with ($\mathsf{ROVEs}$) and without ($\mathsf{ROVE}$) data splitting in training neural networks and solving stochastic programs. Due to space constraints, additional experimental results are provided in Appendix F. In particular, a comprehensive hyperparameter profiling of our algorithms is performed in Appendix F.3 to find empirically well-performing configurations for general use. Unless specified otherwise, all our experiments use the recommended configuration summarized at the end of Appendix F.3.
3.1 Neural Networks for Regression
We consider regression problems with multilayer perceptrons (MLPs) on both synthetic and real data. The base learning algorithm splits the data into training (70%) and validation (30%), and uses Adam to minimize the mean squared error (MSE), with early stopping triggered when the validation improvement falls below 3% between epochs. The architecture details of the MLPs are provided in Appendix F.1. $\mathsf{MoVE}$ is not included in this comparison as it’s applicable to discrete problems only. As a side note, we focus on regression problems since classification models are prevalently trained using the cross-entropy loss which is inherently less prone to heavy-tailed noises thanks to the presence of the logarithm.
Setup for Synthetic Data.    Input-output pairs $(X,Y)$ are generated as $Y=(1/50)\cdot\sum_{j=1}^{50}\log(X_{j}+1)+\varepsilon$, where each $X_{j}$ is drawn independently from $\mathrm{Unif}(0,2+198(j-1)/49)$, and the noise $\varepsilon$ is independent of $X$ with zero mean. We consider both standard Gaussian noise and Pareto noise $\varepsilon=\varepsilon_{1}-\varepsilon_{2}$, where each $\varepsilon_{i}\sim\mathrm{Pareto}(2.1)$. The out-of-sample performance is estimated on a common test set of one million samples. Each algorithm is repeatedly applied to $200$ independently generated datasets to assess the average and tail performance.
Setup for Real Data.    We use six datasets from the UCI Machine Learning Repository (Blake, 1998): Wine Quality (Cortez et al., 2009), Bike Sharing (Fanaee-T, 2013), Online News (Fernandes et al., 2015), Appliances Energy (Candanedo, 2017), Superconductivity (Hamidieh, 2018), and Gas Turbine Emission (gas, 2019). The data is standardized (zero mean, unit variance). To evaluate the average and tail performance, we permute each dataset $100$ times, and each time use the first half for training and the second for testing.
@@ 3.1 Neural Networks for Regression
Results.    As shown in Figure 1, in heavy-tailed noise settings (Figures 1a–1c), both $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ significantly outperform the base algorithm in terms of both expected out-of-sample MSE and tail performance under all sample sizes $n$. Notably, the performance improvement becomes more pronounced with deeper networks ($H=8$), indicating that the benefits of $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ are more apparent in models with higher expressiveness and lower bias.
@@
In light-tailed settings (Figures 1d–1f), $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ show comparable expected out-of-sample performance to the base when $H=4$, but outperform the base as $H$ increases. Additionally, $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ outperform the base in tail probabilities even when $H=4$. This indicates that $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ provide better generalization as the model complexity grows even for light-tailed problems. Similar results for MLPs with $2$ and $6$ hidden layers can be found in Appendix F.4, where results on least squares regression and Ridge regression are also provided.
Figure 2 shows a comparison with bagging that resembles our method most closely among existing ensemble methods as both involve repeated training on randomly drawn subsamples. We implement bagging, or subagging to be precise, on the MLP with $H=4$ hidden layers by averaging the predictions of the repeatedly trained MLPs. The same subsample size and ensemble size are used for our methods and bagging to ensure a fair comparison. Whether bagging or our method wins depends on the tail heaviness: $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ exhibit relatively inferior test performance when the noise has a shape of $2.1$, but outperform bagging as the tail of the noise gets heavier towards a shape of $1.1$.
@@
On real datasets (Figure 3), $\mathsf{ROVE}$ exhibits much lighter tails compared to the base on three datasets, and similar tail behavior on the other three. $\mathsf{ROVEs}$, however, underperforms the base in these real-world scenarios, potentially due to the data split that compromises its statistical power.
@@ 3.2 Stochastic Programs
Setup.    We consider four discrete stochastic programs: resource allocation, supply chain network design, maximum weight matching, and stochastic linear programming, and continuous mean-variance portfolio optimization. All problems are designed to possess heavy-tailed uncertainties. The base learner for all the problems is the SAA. Details of the problems are deferred to Appendix F.2 and results using DRO as the base learner are provided in Appendix F.4.
@@
Results.    Figure 4 shows that our ensemble methods significantly outperform the base algorithm in all cases except for the linear program case (Figure 4(d)). Notably, in the linear program case, $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ still outperform the base, demonstrating their robustness, while $\mathsf{MoVE}$ performs slightly worse than the base under small sample sizes. Comparing $\mathsf{ROVE}$ and $\mathsf{ROVEs}$, $\mathsf{ROVE}$ consistently exhibits superior performance than $\mathsf{ROVEs}$ in all cases.
@@
When there is a unique optimal solution, $\mathsf{MoVE}$ and $\mathsf{ROVE}$ perform similarly, both generally better than $\mathsf{ROVEs}$, as seen in Figures 4(a)-4(c). However, in cases with multiple optima (Figure 4(d)), the performance of $\mathsf{MoVE}$ deteriorates while $\mathsf{ROVE}$ and $\mathsf{ROVEs}$ stay strong. This is in accordance with our discussion on the advantage of $\epsilon$-optimality vote in Section 2.2. Additional results in Appendix F.4 shall further explain that optima multiplicity weakens the base learner for $\mathsf{MoVE}$ in the sense of decreasing the $\eta_{k,\delta}$ and hence inflating the tail bound in Theorem 2.1.
@@
The running time comparison in Figure 4(f) shows that, despite requiring multiple runs on subsamples, our methods do not necessarily incur a higher computation cost compared to running the base algorithm on the full sample, and can even be advantageous under large sample sizes. This is because, in problems like DRO (Ben-Tal et al., 2013; Mohajerin Esfahani & Kuhn, 2018) and two-stage stochastic programming, sampled-based optimization often has a problem size that grows at least linearly with the sample size and requires a superlinearly growing computation cost. Subsampled optimizations, as performed in our approach, are smaller and more manageable. In general, the computation efficiency of our method is ensured by the observation that no more than $\mathcal{O}(n/k)$ subsamples are needed as suggested by our theory and that training on subsamples can be easily parallelized.
@@
Recommended method.    Among the three proposed ensemble methods, $\mathsf{ROVE}$ is the preferred choice over $\mathsf{MoVE}$ and $\mathsf{ROVEs}$ for general use as it’s applicable to both discrete and continuous problems and consistently delivers superior performance across all scenarios.